# Insensitive edge solitons in non-Hermitian topological lattices

### v1

@@ Abstract
In this work, we demonstrate that the synergetic interplay of topology, nonreciprocity and nonlinearity is capable of unprecedented effects. We focus on a nonreciprocal variant of the Su-Shrieffer-Heeger chain with local Kerr nonlinearity. We find a continuous family of non-reciprocal edge solitons (NESs) emerging from the topological edge mode, with near-zero energy, in great contrast from their reciprocal counterparts. Analytical results show that this energy decays exponentially towards zero when increasing the lattice size. Consequently, despite the absence of chiral symmetry within the system, we obtain zero-energy NESs, which are insensitive to growing Kerr nonlinearity. Even more surprising, these zero-energy NESs also persist in the strong nonlinear limit. Our work may enable new avenues for the control of nonlinear topological waves without requiring the addition of complex chiral-preserving nonlinearities.
@@
During the last two decades, great progress has been achieved in understanding topological systems [1, 2, 3, 4, 5, 6, 7, 8, 9]. The promise is the developments of technological devices exploiting localized edge waves immune to fluctuations and defects. The manifestations of these topological edge modes (TMs) were observed in experiments in photonic [4, 10, 11, 12], atomic [13, 14], electronic [15, 16] and phononic [17, 18, 19, 20] devices. Another domain which has been extensively debated recently are linear waves in non-Hermitian (NH) media [21]. These media offer promising solutions in implementing unidirectional, broadband waveguides and amplifiers [22], known to host a plethora of new phenomena like non-Hermitian skin effect (NHSE) [23, 24], complex-frequency and exceptional points [25]. In most cases, the NHSE originates from the asymmetry between the couplings which constraints most of the system normal modes (NMs) to be localized on a single interface of the system. This phenomenon has been experimentally demonstrated in photonics [26, 27], electronics [28], acoustics [29, 30] and mechanics [31, 32, 33]. Owing to this, it was shown that NHSE and TM can interact, leading to unprecedented level of manipulation of the latter, e.g., in phononic [32, 34, 35, 33] and electric [36] systems.
Interestingly, the inclusion of nonlinear features in NH topological insulators has recently seen growing interest. Exciting developments like nonlinear NHSE [37, 38, 39, 40], nonlinearly induced NH phase transitions [41] and nonlinear wave acceleration [42] have already been demonstrated. Furthermore, experiments featuring nonreciprocity revealed the realization of lossless unidirectional waveguides [43, 44, 45, 42]. Following these efforts, we raise the question of what is the fate of nonlinear topological solitons (or breathers) [46, 47, 48, 49, 50, 51, 52, 53, 54] in the presence of nonreciprocity? Assuming that nonlinearity is perturbative, we can prematurely collect some elements of answers. Indeed, it was demonstrated [55, 56] that NH topological systems are exponentially sensitive to perturbations at the boundaries.
Figure 1: (a): The nonreciprocal SSH model considered in this work with Kerr nonlinearity. (b) Phase diagram $(\gamma,s)$ of the linearized model. The insets indicate the localization of the right TM and SMs.
@@
Our results are based on finite lattices whose governing equations read
@@
where $E$ is the energy, $\sigma=\pm 1$ the nonlinear coefficient, $s$ and $t\pm\gamma$ the positive nearest-neighbour couplings, and $\gamma>0$ quantifies the strength of nonreciprocity. Here $u_{a,n}$ is the amplitude at cell with index $n$ and sublattice $a=\{A,B\}$. In the linear limit, $\lvert u_{a,n}\rvert^{2}\rightarrow 0$, Eq (2), reduces to a linear eigenvalue problem $H\vec{u}=E\vec{u}$, with the $H$ being a non-Hermitian and chiral matrix [58]. Hereafter, we consider a chain of $N$ cells with $12N-$ 1 sites and open boundary conditions. The chirality ensures all bulk eigenenergies to come in pairs $E_{k}=\pm\sqrt{s^{2}+\tilde{t}^{2}+2s\tilde{t}\cos\left(\frac{k\pi}{2N}\right)}$ with $\lvert k\rvert=1,2,\ldots,N-1$ and $\tilde{t}=\sqrt{t^{2}-\gamma^{2}}$. Regarding their associated NMs, the non-Hermiticity implies the $k$-th right and left NMs, respectively $\vec{u}_{k}$ and $\vec{v}_{k}$ to be in general different [59]. Furthermore, these NMs are skin modes (SMs) satisfying $u_{k,a,n}\sim\beta^{(n-1)/2}$ and $v_{k,a,n}\sim\beta^{-(n-1)/2}$ with $\beta=(t+\gamma)/(t-\gamma)$ quantifying their localization [58]. As such, the right SMs are always localized on the left boundary, while the left ones are localized on the right sides of the chain, Fig. 1(b).
Figure 2: The numerically obtained $E$ of the NES against $I$, (a) for regions (I) and (III) and (b) for (II) with $\gamma=0.3$, $0.825$ and $0.6$ respectively. The dashed lines in panel (b) show the corresponding Hermitian SSH model $(\gamma=0)$ for comparison [60]. The insets depict the localization sides of the TMs.
In addition, the choice of $12N-$ 1 sites and chiral symmetry guarantee the presence of a NM with $E_{0}=0$, pinned as TM. It has zero support on the B-sublattice, $u_{0,B,n}=v_{0,B,n}=0$, while the A-sublattice $A$ satisfies
@@
where $r_{R}=-s/(t-\gamma)$ and $r_{L}=-s/(t+\gamma)$. The localization of the TM is distinct from the SMs and can be used to identify three different regions, namely:
$\mathrm{(I)}:$ $s+\gamma<t,\quad$ $\lvert r_{R}\rvert,\lvert r_{L}\rvert<1,$
$\mathrm{(II)}:$ $\sqrt{s^{2}+\gamma^{2}}<t<s+\gamma,\quad$ $\lvert r_{R}\rvert>1,\lvert r_{L}\rvert<1,$ (4)
$\mathrm{(III)}:$ $\sqrt{s^{2}+\gamma^{2}}>t,\quad$ $\lvert r_{R}\rvert,\lvert r_{L}\rvert>1.$
as highlighted in Fig. 1(b). In region (I) both the right and left TMs, are localised on the opposite side compared to right SMs. On the other hand, for region (III) the nonreciprocity is dominating. It follows that the right and left TMs are opposite, and on the same side as their respective SMs with similar localization. Finally region (II) looks like (III) with the difference being the TMs and SMs exhibit distinct localization (see also Ref. [61]).
@@
Figure 3: (a) Dependence of the energy of the NESs at $I=0.5$ against $\gamma$ with $s=0.6$. The black lines are the spectrum of the linearized limit. (b) Representative cases of NESs for the Hermitian ($\gamma=0$) and NH SSH chains with $\gamma=0.3$, $0.5$ and $0.6$ and $0.825$ respectively the star, triangle, circle, square and diamond dots. (c) Energy sensitivity of the NESs as function of the lattice size for (I) $\gamma=0.3$, (II) $0.6$ and (III) $0.825$.
@@
We now numerically calculate the nonlinear solutions of Eqs. (2) stemming from the TM, using a pseudo-arclength solver [62, 63, 64] and follow families of solutions of increasing intensity, i.e. $I=\sum_{l}\lvert u_{l}\rvert^{2}$. The solver returns both the $E$ and the $u_{a,n}$ of the obtained solutions. Figure 2(a) shows the dependence of $E$ as a function of $I$ for representative cases of the found families of NESs on a chain of $N=17$ with $\gamma=0.3$ and $0.825$ ($s=0.6$) in regions (I) and (III) respectively. It depicts a behavior similar to the Hermitian SSH model with $\gamma=0$ [see dashed curve in Fig. 2(b)]. That is to say, the energy departs from the origin, growing with increasing intensity till approaching the linear upper band for $\sigma>0$. This behaviors is associated to the chiral symmetry breaking induced by Kerr nonlinearity which do not guarantee the presence of modes at $E=0$.
@@
An unexpected result is found in region (II). Indeed, despites the aforementioned chiral symmetry breaking we find the family above has $E\approx 0$ when increasing the amplitude, as shown in case of $\gamma=0.6$ and $s=0.6$ in Fig. 2(b). In order to explain the origin of this behavior, we rely on the perturbation theory, assuming small amplitude nonlinear solutions originating from the $k$-th NM expand as $\vec{u}_{k}=\vec{u}_{k}^{0}+\epsilon\vec{u}_{k}^{1}+\mathcal{O}(\epsilon^{2})$ with the energy given by $E_{k}=E_{k}^{0}+\epsilon E_{k}^{1}+\mathcal{O}(\epsilon^{2})$. Here the $\epsilon$ controls the amplitude and the $(\vec{u}_{k}^{0},E_{k}^{0})$ are retrieved resolving the linearized limit. The details of the first order perturbation analysis $(\vec{u}_{k}^{1},E_{k}^{1})$ can be found in the supplementary [58]. Consequently, we define the rate of change of the energy to the intensity refereed to as energy sensitivity factor,
@@
$\mathcal{S}_{E}^{k}=\frac{\partial E_{k}}{\partial I}\approx\rho_{k}E_{k}^{1}(\sigma),$ (5)
where $\rho_{k}^{-1}=\vec{u}_{k}^{0,T}\vec{u}_{k}^{0}$ is analogous to the Petermann factor [65]. One of the main results of our work is the following relation unique to the NES:
@@
with $S_{0}$ being a positive nonzero constant [58]. This expression implies that the energy shift of the NES is obtained solely by examining the right and left TMs. Thus, using Eqs. (4) and (6), we can predict behavior of the families of NESs emerging from the TM within the parameter space of the linearized model, Fig. 1(b). More specifically; for sufficiently large $N$ values; we find that in regions (I) and (III) the $\mathcal{S}^{0}_{E}\sim 1$. Thus a similar energy grows against the intensity must be observed for all $N$, confirming the results of Fig. 2(a). On the other hand, in region (II) we obtain $\mathcal{S}^{0}_{E}\sim\left(r_{L}r_{R}\right)^{N}$. It follows rewriting $r_{L}r_{R}=e^{-\alpha_{E}}$ with $10<\alpha_{E}(r_{R},r_{L})<$ 1, that the energy sensitivity is exponentially decaying with $N$. As such, for large lattices we obtain zero-energy NESs, explaining the unexpected numerical observation of Fig. 2(b).
Figure 4: (a) Projection, $C_{k}$, of the NESs in Fig. 3(b) into the NMs. The $\lvert C_{k}\rvert$ are rescaled by their maximum for clarity. (b) Similar to Fig. 3(c) but for the wavefunction’s sensitivity. (c) Stability of the NES at fixed $I=0.5$, across the phase diagram of Fig. 1(b). We display the largest eigenvalue, $\lambda$, of the stability matrix [58]. The white line and curve represent $s=t-\gamma$ and $s^{2}=t^{2}-\gamma^{2}$ respectively. Inset: Dependence of $\lambda$ against $\gamma$ at $s=0.6$ (see dashed line).
@@
We scan the system’s parameter space in Fig. 1(b), performing numerical continuation on a chain of $N=17$ at fixed nonlinearity, $I=0.5$, varying the $\gamma$ along the section $s=0.6$. Along this line, the $\gamma=0.4$ delimits regions (I) and (II) and $\gamma=0.8$, regions (II) and (III). The results are shown in Fig. 3(a). Clearly at $\gamma=0$ in the Hermitian case, an energy shift away from $E=0$ is seen [star dot]. As $\gamma\rightarrow 0.4$ this energy shift remains non-trivial, while tending to decrease with increasing values of $\gamma$ across region (I). Representative NESs of this region are also shown in Fig. 3(b) for $\gamma=0$ and $0.3$ respectively the star and triangle dots. These NESs stay localized at the same side as their TM, yet developing support also on the $B$-sublattice. In addition, similar results are also seen for region (III) where $\gamma>0.8$ [Fig. 3(a)], with a representative NES depicted by the diamond dot in Fig. 3(b). These observations, once again, are due to the chiral symmetry breaking owning to the Kerr nonlinearity.
Bearing the above in mind, entering region (II) with $0.80.4<\gamma<$ 0.8, we find this family of NESs at $I=0.5$ has $E\approx 0$, Fig. 3(a). In addition, we depict representative NESs with $\gamma=0.5$ and $0.6$ respectively the circle and square dots in Fig. 3(b). Their profiles clearly show localized shapes, with support also on the B-sublatttice.
@@
Figure 3(c) displays the dependence of the sensitivity factor, $\mathcal{S}_{E}^{0}$ against the lattice size $N$. The measures are obtained as in Fig. 3(a) for representative parameter sets (I) $\gamma=0.3$, (II) $0.6$ and (III) $0.825$ with $s=0.6$. For regions (I) and (III), a clear saturation of the $\mathcal{S}_{E}^{0}$ to values in the interval $[0.1,1]$ is seen. On the other hand, for $\gamma=0.6$ in region (II), we find an exponentially decaying sensitivity factor, $\mathcal{S}_{E}^{0}\sim e^{-0.6N}$. It is worth emphasizing that we have also checked that this insensitivity of the NESs does not holds for the nonlinear SMs and are robust whilst the addition of disorder [58].
@@
Let us now focus on the shape of the NES rather than its energy. The perturbation theory can also be used to define the wavefunction sensitivity
$\mathcal{S}_{\vec{u}}^{k}=\frac{\partial\vec{u}_{k}}{\partial I}\approx\rho_{k}V_{k}^{1}(\sigma),$ (7)
@@
with $V_{k}^{1}=\lvert\sum_{l\neq k}C_{l}^{1}(\sigma)u_{l}^{0}\rvert$ being a mode overlap integral [58]. Consequently, the $\mathcal{S}_{\vec{u}}^{0}$ quantifies how strong the TM couples to the SMs when increasing the intensity [58]. We project the NESs of Fig. 3(b) onto the NM basis, $\vec{u}=\sum_{k}C_{k}\vec{u}_{k}^{0}$. The $\lvert C_{k}\rvert$, are shown in Fig. 4(a). It follows that despite these NESs being obtained for the same nonlinearity, $I=0.5$, the TM strongly couples with the SMs [star, rectangle and diamond dots] in regions (I) and (III), while in region (II), this coupling is smaller [circle and square dots]. Further, the dependence of $\mathcal{S}_{\vec{u}}^{0}$ as function of $N$ shows that this mode couplings growth or remains constant with increasing $N$ in regions (I) and (III) respectively. Indeed, we expect that as the lattice size grows, the number of NMs the TM pairs with increases. Remarkably, region (II), evades this anticipation and the $\mathcal{S}_{\vec{u}}^{0}$ tends to exponentially vanish. In the supplementary [58], we shows that these results are independent of $I$. We also demonstrate that, similar to the TM case, the shape of finite amplitude NESs can be modified by tuning $\gamma$ (see also [33]).
@@
With respect to the linear stability of the NESs, the latter is obtained monitoring a small deviation $\chi w_{a,n}(\tau=0)$ from the NES $u_{a,n}$ of the form $\psi_{a,n}=\left(u_{a,n}+\chi w_{a,n}\right)e^{iE\tau}$ with $\chi\ll 1$ and $\tau$ being the temporal variable [66]. Considering $w_{a,n}\sim e^{i\lambda\tau}$ leads to a linear eigenvalue problem $\lambda\vec{w}=Z\vec{w}$ [58]. Thus, instabilities are signaled by the imaginary part of $\lambda$. The largest value of the imaginary part of $\lambda$ is plotted in Fig. 4(c), for the parameter space of Fig. 1(b) using the same setting as in Fig. 3. We observe two regions of stable NESs [blue islands]. The stability in the lower islands, is likely due to large band gap, as $s\rightarrow 0$. On the other hand, despite region (II), being in the neighbourhood of the band closing the obtained NESs are in general linearly stable, see also inset of Fig. 4(c).
@@
Figure 5: The energy, $E$ against intensity, $I$ for the families of nonlinear modes emerging from the NMs [we show $8$ of them] on a chain of $N=17$ cells with (II) $\gamma=0.6$ and $s=0.6$. The insets (left) zoom into a transition and (right) depict the projection of the NES of $E\approx 0$ at $I=2.75$ into the NMs basis (red dots).

### v2

# Insensitive edge solitons in a non-Hermitian topological lattice

@@ Abstract
In this work, we demonstrate that the synergetic interplay of topology, nonreciprocity and nonlinearity is capable of unprecedented effects. We focus on a nonreciprocal variant of the Su-Shrieffer-Heeger chain with local Kerr nonlinearity. We find a continuous family of non-reciprocal edge solitons (NES) emerging from the topological edge mode, with near-zero energy, in great contrast from their reciprocal counterparts. Analytical results show that this energy decays exponentially towards zero when increasing the lattice size. Consequently, despite the absence of chiral and sublattice symmetries within the system, we obtain zero-energy NES, which are insensitive to growing Kerr nonlinearity. Even more surprising, these zero-energy NES also persist in the strong nonlinear limit. Our work may enable new avenues for the control of nonlinear topological waves without requiring the addition of complex chiral- or sublattice-preserving nonlinearities.
@@
During the last two decades, great progress has been achieved in understanding topological systems [1, 2, 3, 4, 5, 6, 7, 8, 9]. The promise is the developments of technological devices exploiting localized edge waves immune to fluctuations and defects. The manifestations of these topological edge modes were observed in experiments in photonic [4, 10, 11, 12], atomic [13, 14], electronic [15, 16] and phononic [17, 18, 19, 20] devices. Another domain which has been extensively debated recently are linear waves in non-Hermitian (NH) media [21]. These media offer promising solutions in implementing unidirectional, broadband waveguides and amplifiers [22], known to host a plethora of new phenomena like non-Hermitian skin effect (NHSE) [23, 24, 25], complex-frequency and exceptional points [26]. In most cases, the NHSE originates from the asymmetry between the couplings which constraints most of the system normal modes (eigenmodes) to be localized on a single interface of the system. This phenomenon has been experimentally demonstrated in photonics [27, 28], electronics [29], acoustics [30, 31] and mechanics [32, 33, 34]. Owing to this, it was shown that NHSE and topological modes can interact, leading to unprecedented level of manipulation of the latter, e.g., in phononic [33, 35, 36, 34] and electric [37] systems.
Interestingly, the inclusion of nonlinear features in NH topological insulators has recently seen growing interest. Exciting developments like nonlinear NHSE [38, 39, 40, 41, 42, 43], nonlinearly induced NH phase transitions [44] and nonlinear wave acceleration [45] have already been demonstrated. Furthermore, experiments featuring nonreciprocity revealed the realization of lossless unidirectional waveguides [46, 47, 48, 45]. Following these efforts, we raise the question of what is the fate of nonlinear topological solitons (or breathers) [49, 50, 51, 52, 53, 54, 55, 56, 57] in the presence of nonreciprocity? Assuming that nonlinearity is perturbative, we can prematurely collect some elements of answers. Indeed, it was demonstrated [58, 59] that NH topological systems are exponentially sensitive to perturbations at the boundaries.
Figure 1: (a): The nonreciprocal SSH model considered in this work with Kerr nonlinearity. (b) Phase diagram $(\gamma,s)$ summarizing the localization characteristics of the TM right and left eigenvectors. Representative cases are shown for $s=0.6t$ with (I) $\gamma=0.3t$, (II) $\gamma=0.6t$, and (III) $\gamma=0.825t$ (see black circles). Right: The localization behavior of the right and left eigenvectors of the TM in regions (I), (II), and (III).
@@
Our results are based on finite lattices whose governing equations read [61, 62]
@@
with $E$ being the energy, $\sigma=\pm 1$ the nonlinear coefficient, $s$ and $t\pm\gamma$ the positive nearest-neighbour couplings, and $\gamma>0$ quantifying the strength of nonreciprocity.
Figure 2: The numerically obtained $E$ of the NES as a function of $I$ for representative cases at $s=0.6$ is shown in (a) for regions (I) and (III) with $\gamma=0.3$ [blue curve] and $\gamma=0.825$ [gray curve]. (b) Same as in (a) but for region (II) with $\gamma=0.6$ [orange curve], see also Fig. 1(b). The dashed line in (b) represents the corresponding result for the Hermitian SSH model with $\gamma=0$, for comparison. The insets depict the localization characteristics of the right and left eigenvectors of the TM.
Here $u_{a,n}$ is the amplitude at cell with index $n$ and sublattice $a=\{A,B\}$, Fig. 1(a). In the linear limit, $\lvert u_{a,n}\rvert^{2}\rightarrow 0$, Eq (2), reduces to a linear eigenvalue problem $H\vec{u}=E\vec{u}$, with the $H$ being a non-Hermitian matrix possessing chiral and sublattice symmetries (CS and SLS) [63, 64, 65, 66, 61, 62]. Hereafter, we consider a chain of $N$ cells with $12N-$ 1 sites and open boundary condition (OBC), Fig. 1(a).
The SLS/CS ensures all bulk eigenenergies to come in pairs $E_{l}=\pm\sqrt{s^{2}+\tilde{t}^{2}+2s\tilde{t}\cos\left(\frac{l\pi}{2N}\right)}$ with $\lvert l\rvert=1,2,\ldots,N-1$, labeling the eigenmodes energies and $\tilde{t}=\sqrt{t^{2}-\gamma^{2}}$. Regarding their associated $22N-$ 2 eigenvectors, non-Hermiticity implies the $l$-th right ($\vec{u}_{l}$) and left ($\vec{v}_{l}$) eigenvectors are in general different [67]. These eigenvectors exhibit a localised (skinny) profile where each site amplitude satisfy
$u_{l,a,n}\sim d_{R}^{n-1}\mbox{ and }v_{l,a,n}\sim d_{L}^{n-1},$ (3)
with $d_{R}=d_{L}^{-1}=\sqrt{(t+\gamma)/(t-\gamma)}$ (see supplementary [62]). Thus the right and left eigenvectors of bulk modes have equivalent localization, but are respectively located on the left and right boundaries of the system. This localization characteristics, $d_{R}d_{L}=1$, is independent of the parameter coordinates in Fig. 1(b). In addition, the extent of localization solely depends on the non-reciprocal strength, $\gamma$, similar to the Hatano-Nelson model. These bulk modes are commonly referred to as skin modes.
In addition to the bulk modes, the choice of $12N-$ 1 sites and SLS/CS guarantee the presence of a zero-energy mode with index $l=0$, i.e. $E_{0}=0$, pinned as topological mode (TM). Its left and right eigenvectors have zero support on the B-sublattice, $u_{0,B,n}=v_{0,B,n}=0$, while the A-sublattice satisfies
@@
where $r_{R}=-s/(t-\gamma)$ and $r_{L}=-s/(t+\gamma)$. Importntly, the localization ratios $r_{R,L}$ of the TM behave very differently than those of the bulk modes and in general $r_{R}r_{L}\neq 1$. In this context, we can identify three different regions for the TM, namely:
$\mathrm{(I)}:s+\gamma<t,\quad\lvert r_{R}\rvert<1,\lvert r_{L}\rvert<1,$
$\mathrm{(II)}:\sqrt{s^{2}+\gamma^{2}}<t<s+\gamma,\quad\lvert r_{R}\rvert>1,\lvert r_{L}\rvert<\lvert r_{R}\rvert^{-1}<1,$
$\mathrm{(III)}:\sqrt{s^{2}+\gamma^{2}}>t,\quad\lvert r_{R}\rvert>1,1>\lvert r_{L}\rvert>\lvert r_{R}\rvert^{-1},$ (5)
as highlighted in Fig. 1(b). In region (I) both the right and left eigenvectors of the TM, are located on the same side. On the other hand, in regions (II) and (III), non-reciprocity dominates, causing these eigenvectors to be located opposite to each other. Consequently, we examine their localization: in region (II), the right eigenvector of the TM is less localized compared to its left counterpart, and vice versa in region (III). From here on, we set $t=1$ without loss of generality.
@@
Figure 3: (a) Dependence of the energy of the NES at $I=0.5$ against $\gamma$ with $s=0.6$. The black lines are the spectrum of the linear limit. (b) Profiles of some representative cases of the NES [see panel (a)]. (c) Energy sensitivity of the NES as function of the lattice size for (I) $\gamma=0.3$, (II) $0.6$ and (III) $0.825$.
@@
We now numerically calculate the nonlinear solutions of Eq. (2) stemming from the TM, using a pseudo-arclength solver [68, 69, 70] and follow families of solutions of increasing intensity, i.e. $I=\sum_{j}\lvert u_{j}\rvert^{2}$ ($j=1,2,\ldots,2N-1$) using $\sigma=1$ [71]. The solver returns both the $E$ and the $u_{a,n}$ of the obtained solutions. Figure 2(a) shows the dependence of $E$ as a function of $I$ for representative cases of the families of NES on a chain of $N=17$ with $\gamma=0.3$ and $0.825$ ($s=0.6$) in regions (I) and (III) respectively. It depicts a behavior similar to the Hermitian SSH model with $\gamma=0$ [see dashed curve in Fig. 2(b)]. That is to say, the energy departs from the origin, growing with increasing intensity till approaching the upper band of the linear spectrum when $\sigma>0$, indicating a significant effect of nonlinearity. This behavior is associated to the SLS/CS breaking induced by Kerr nonlinearity which does not guarantee the presence of solutions at $E=0$.
@@
An unexpected result is found in region (II). Indeed, despite the aforementioned SLS/CS breaking, we find the family above has $E\approx 0$ when increasing the amplitude, as shown in case of $\gamma=0.6$ and $s=0.6$ in Fig. 2(b). In order to explain the origin of this behavior, we rely on the perturbation theory, assuming small amplitude nonlinear solutions originating from the $l$-th eigenmode expand as $\epsilon^{-1/2}\vec{u}_{l}=\vec{u}_{l}^{(0)}+\epsilon\vec{u}_{l}^{(1)}+\mathcal{O}(\epsilon^{2})$ with the energy given by $E_{l}=E_{l}^{(0)}+\epsilon E_{l}^{(1)}+\mathcal{O}(\epsilon^{2})$. Here the $\epsilon$ is the perturbation parameter which controls the amplitude and the $(\vec{u}_{l}^{(0)},E_{l}^{(0)})$ are retrieved resolving system in its linearized limit. The details of the first order perturbation analysis $(\vec{u}_{l}^{(1)},E_{l}^{(1)})$ can be found in the supplementary [62]. Consequently, we define the rate of change of the energy to the intensity referred to as energy sensitivity factor,
@@
$\mathcal{S}_{E}^{l}=\frac{\partial E_{l}}{\partial I}\approx\frac{\vec{v}_{l}^{(0)T}\Gamma_{l}(\sigma)\vec{u}_{l}^{(0)}}{\left(\vec{u}_{l}^{(0)T}\vec{u}_{l}^{(0)}\right)\left(\vec{v}_{l}^{(0)T}\vec{u}_{l}^{(0)}\right)}=\rho_{l}E_{l}^{(1)}(\sigma),$ (6)
where $\rho_{l}^{-1}=\vec{u}_{l}^{(0)T}\vec{u}_{l}^{(0)}$ and $\{^{T}\}$ stands for the vector transpose, see supplemental [62]. In addition, the $\left[\Gamma_{l}\right]_{m,j}=\sigma\left\lvert u_{l,m}^{(0)}\right\rvert^{2}\delta_{m,j}$ is the diagonal matrix accounting for the nonlinear perturbation. According to Eq.(6), we see that the energy shift stems from two contributions: the nonlinearity through $\delta_{l}\sim\vec{v}_{l}^{(0)}\Gamma_{l}(\sigma)\vec{u}_{l}^{(0)}$, and the non-orthogonality of the eigenvectors of the underlying linear system quantified by the Petermann-like factor $K_{l}^{-1}\sim(\vec{u}_{l}^{(0)T}\vec{u}_{l}^{(0)})(\vec{u}_{l}^{(0)T}\vec{v}_{l}^{(0)})$.
Figure 4: (a) Projection, $C_{l}$, of the NES of Fig. 3(b) into the eigenmode basis. The $\lvert C_{l}\rvert$ are rescaled by their maximum for clarity. (b) Similar to Fig. 3(c) but for the wavefunction’s sensitivity. (c) Linear stability of the NES at fixed $I=0.5$, across the phase diagram of Fig. 1(b). We display the largest imaginary part of $\lambda$. The green solid curves $s=t-\gamma$ and $s^{2}=t^{2}-\gamma^{2}$, delimiting the three regions characterizing the TM right and left eigenvectors. Right: Dependence of $\max\{\rm{Im}\lambda\}$ against $\gamma$ at $s=0.6$ (see dashed line).
One of the main results of our work is the following relation specific to the NES:
@@
with $S_{0}$ being a positive nonzero constant [62]. This relation suggests that the imbalance between the right and left eigenvectors, which is unique to the TM, i.e. $r_{R}r_{L}\neq 1$, can drastically change the nonlinear energy shift [Eq. (7)] under weak nonlinearity. Indeed as the $r_{R}$ and $r_{L}$ vary within the parameter space of Fig. 1(b), this expression implies that the physical mechanism at the origin of the energy shift of the TM results from the interplay between nonlinearity, $\delta_{0}\sim(r_{L}r_{R}r_{R}r_{R})^{N}-1$ and non-orthogonality, $K_{0}^{-1}\sim[(r_{R}r_{R})^{N}-1][(r_{L}r_{R})^{N}-1]$. More specifically, using Eq.(7) and looking at large values of $N$, we explicitly get that in regions (I) and (III) the $\mathcal{S}^{0}_{E}\sim 1$. Thus energy is expected to grow with intensity confirming the results of Fig. 2(a). On the other hand, in region (II) we obtain $\mathcal{S}^{0}_{E}\sim\left(r_{L}r_{R}\right)^{N}\equiv e^{-\alpha_{E}N}$ with $10<\alpha_{E}(r_{R},r_{L})<$ 1, leading to an energy sensitivity which exponentially decays with $N$. As such, for large lattices we obtain zero-energy NES, explaining the unexpected numerical observation of Fig. 2(b).
@@
Another way to illustrate this result is obtained by scanning the system’s parameter space of Fig. 1(b), showing the energy of the NES of fixed nonlinearity, $I=0.5$, varying $\gamma$ along the section with $s=0.6$. Along this line, the $\gamma=0.4$ delimits regions (I) and (II) and $\gamma=0.8$, regions (II) and (III). The result is shown in Fig. 3(a) with the blue dots. Clearly at $\gamma=0$ in the Hermitian case, an energy shift away from $E=0$ is seen [star]. As $\gamma\rightarrow 0.4$ this energy shift remains non-trivial, while tending to decrease with increasing values of $\gamma$ across region (I). Representative NES of this region are also shown in Fig. 3(b) for $\gamma=0$ and $0.3$ respectively the star and triangle dots. These NES stay localized at the same side as their linear counterparts, yet developing support also on the $B$-sublattice. In addition, similar results are also seen for region (III) where $\gamma>0.8$ [Fig. 3(a)], with a representative NES depicted by the diamond dot in Fig. 3(b). These observations, once again, are due to the SLS/CS breaking owning to the Kerr nonlinearity.
@@
Bearing the above in mind, entering region (II) with $0.80.4<\gamma<$ 0.8, we find this family of NES at $I=0.5$ has $E\approx 0$, Fig. 3(a). In addition, we depict representative NES with $\gamma=0.5$ and $0.6$ respectively the circle and square dots in Fig. 3(b). Their profiles clearly show localized shapes, with support also on the B-sublatttice. Figure 3(c) displays the dependence of the sensitivity factor, $\mathcal{S}_{E}^{0}$ against the lattice size $N$. The measures are obtained as in Fig. 3(a) for representative parameter sets (I) $\gamma=0.3$, (II) $0.6$ and (III) $0.825$ with $s=0.6$. For regions (I) and (III), a clear saturation of the $\mathcal{S}_{E}^{0}$ to values in the interval $[0.1,1]$ is seen. On the other hand, for $\gamma=0.6$ in region (II), we find an exponentially decaying sensitivity factor, $\mathcal{S}_{E}^{0}\sim e^{-0.6N}$. It is worth emphasizing that we have also checked that this insensitivity of the NES does not hold for the nonlinear modes emerging from the bulk and are robust whilst the addition of disorder [62].
@@
Let us now focus on the shape of the NES rather than their energy. The perturbation theory can also be used to define the wavefunction sensitivity
$\mathcal{S}_{\vec{u}}^{l}=\frac{\partial\vec{u}_{l}}{\partial I}\approx\rho_{l}V_{l}^{(1)}(\sigma),$ (8)
@@
with $V_{l}^{(1)}=\lvert\sum_{m\neq l}c_{l,m}(\sigma)u_{m}^{(0)}\rvert$ being a mode overlap integral [62]. Consequently, the $\mathcal{S}_{\vec{u}}^{0}$ quantifies how strong the TM couples to the bulk modes when increasing the intensity [62]. We project the NES of Fig. 3(b) onto the eigenmode basis, $\vec{u}=\sum_{l}C_{l}\vec{u}_{l}^{(0)}$ and the corresponding coefficients $\lvert C_{l}\rvert$, are shown in Fig. 4(a) for fixed $s$, varying the values of $\gamma$ spanning across regions (I), (II) and (III). It follows that despite these NES being obtained for the same nonlinearity, $I=0.5$, the TM strongly couples to the bulk [star, triangle and diamond dots] in regions (I) and (III), while in region (II), this coupling is smaller [circle and square dots]. Further, the dependence of $\mathcal{S}_{\vec{u}}^{0}$ as function of $N$ in Fig. 4(b) shows that this mode coupling grows or remains constant with increasing $N$ in regions (I) and (III) respectively. Indeed, we expect that as the lattice size grows, the number of eigenmodes, the TM pairs with, increases. Remarkably, region (II), evades this anticipation and the $\mathcal{S}_{\vec{u}}^{0}$ tends to exponentially vanish. It follows that the TM do not practically couples with the bulk modes in large lattices. In the supplementary [62], we show that these results are independent of $I$. We also demonstrate that, similar to the TM case, the shape of finite amplitude NES can be modified by tuning $\gamma$.
@@
To complete our study we move to the linear stability of the NES. The latter is obtained monitoring the time evolution of a small deviation from a nonlinear stationary solution in the form $\psi_{a,n}(\tau)=\left(u_{a,n}+\chi w_{a,n}(\tau)\right)e^{iE\tau}$ with $\chi\ll 1$, $w_{a,n}\propto 1$ and $\tau$ being the temporal variable. After linearization, we obtain a linear evolution equation for the perturbation $w_{a,n}$. Considering $w_{a,n}(\tau)\sim e^{i\lambda\tau}$ leads to a linear eigenvalue problem $\lambda\vec{w}=Z\vec{w}$ [62, 72]. Thus, instabilities are signaled by the imaginary part of $\lambda$. The largest value of the imaginary part of $\lambda$ is plotted as a colormap in Fig. 4(c), for values of the parameters of the system throughout Fig. 1(b) using the same setting as in Fig. 3. We observe two regions of linearly stable NES [blue islands]. The stability in the lower islands, is likely due to large band gap, as $s\rightarrow 0$ and the NES being far away from the linear energy bands. On the other hand, despite region (II), being in the neighbourhood of the band closing, the obtained NES are in general linearly stable, right panel of Fig. 4(c).
@@
Figure 5: The energy, $E$ against intensity, $I$ for the families of nonlinear modes emerging from all eigenmodes [we show $8$ of them] on a chain of $N=17$ cells with (II) $\gamma=0.6$ and $s=0.6$. The insets (left) zoom into a transition and (right) depict the projection of the NES of $E\approx 0$ at $I=2.75$ into the eigenmodes basis (red dots).